Some one please explain!!!
Assume that the population of heights of female college students is approximately normally distributed with mean of 68 inches and standard deviation of 2.75 inches. Show all work.
(A) Find the proportion of female college students whose height is greater than 69 inches.
(B) Find the proportion of female college students whose height is no more than 69 inches.

Answer by stanbon(57967)

Assume that the population of heights of female college students is approximately normally distributed with mean of 68 inches and standard deviation of 2.75 inches. Show all work.
(A) Find the proportion of female college students whose height is greater than 69 inches.
---
z(69) = (69-68)/2.75 = 0.3640
P(x > 69) = P(z > 0.3640) = 0.3581
---------
(B) Find the proportion of female college students whose height is no more than 69 inches.
Ans: 1-0.3581 = 0.6419
===========================
Cheers,
Stan H.
=========